### v1

# Light-A-Video: Training-free Video Relighting via Progressive Light Fusion

Abstract
Recent advancements in image relighting models, driven by large-scale datasets and pre-trained diffusion models, have enabled the imposition of consistent lighting. However, video relighting still lags, primarily due to the excessive training costs and the scarcity of diverse, high-quality video relighting datasets. A simple application of image relighting models on a frame-by-frame basis leads to several issues: lighting source inconsistency and relighted appearance inconsistency, resulting in flickers in the generated videos. In this work, we propose Light-A-Video, a training-free approach to achieve temporally smooth video relighting. Adapted from image relighting models, Light-A-Video introduces two key techniques to enhance lighting consistency. First, we design a Consistent Light Attention (CLA) module, which enhances cross-frame interactions within the self-attention layers to stabilize the generation of the background lighting source. Second, leveraging the physical principle of light transport independence, we apply linear blending between the source video’s appearance and the relighted appearance, using a Progressive Light Fusion (PLF) strategy to ensure smooth temporal transitions in illumination. Experiments show that Light-A-Video improves the temporal consistency of relighted video while maintaining the image quality, ensuring coherent lighting transitions across frames. Project page: https://bujiazi.github.io/light-a-video.github.io/.
Figure 1: Training-free video illumination control. Equipped with a pretrained image relighting model (e.g., IC-Light [63]) and a video diffusion model (e.g., CogVideoX [60] and AnimateDiff [13]), Light-A-Video enables training-free and zero-shot illumination control of any given video sequences or foreground sequences.
00footnotetext: * indicates equal contribution, † indicates corresponding author
1 Introduction
Illumination plays a crucial role in shaping our perception of visual content, impacting both its aesthetic quality and human interpretation of scenes. Relighting tasks [48, 33, 35, 66, 45, 18, 55, 67, 25], which focus on adjusting lighting conditions in 2D and 3D visual content, have long been a key area of research in computer graphics due to their broad practical applications, such as film production, gaming, and virtual environments. Traditional image relighting methods rely on physical illumination models, which struggle with accurately estimating real-world lighting and materials.
To overcome these limitations, data-driven approaches [10, 40, 23, 65, 61, 20] have emerged, leveraging large-scale, diverse relighting datasets combined with pre-trained diffusion models. As the state-of-the-art image relighting model, IC-Light [63] modifies only the illumination of an image while maintaining its albedo unchanged. Based on the physical principle of light transport independence, IC-Light allows for controlled and stable illumination editing, such as adjusting lighting effects and simulating complex lighting scenarios. However, video relighting is significantly more challenging due to maintaining temporal consistency across frames. The scarcity of video lighting datasets and the cost of training further complicate the task. Thus, existing video relighting methods [65] struggle to deliver consistently high-quality results or are limited to specific domains, such as portraits.
Figure 2: Comparison of relighted frames using IC-Light and CLA module frame-by-frame. The line chart depicts the average optical flow intensity per frame. IC-Light shows a noticeable jitter between frames. Conversely, our CLA module achieves more stable light source generation over time, making the motion trend more consistent with the source video.
In this work, we propose a training-free approach for video relighting, named Light-A-Video, which enables the generation of smooth, high-quality relighted videos without any training or optimization. As shown in Fig. 1, given a text prompt that provides a general description of the video and specified illumination conditions, our Light-A-Video pipeline can relight the input video in a zero-shot manner, fully leveraging the relighting capabilities of image-based models and the motion priors of the video diffusion model. To achieve this goal, we initially apply an image-relighting model to video-relighting tasks on a frame-by-frame basis, and observe that the generated lighting source is unstable across frames in the video. This instability leads to inconsistencies in the relighting of the objects’ appearances and significant flickering across frames. To stabilize the generated lighting source and ensure consistent results, we design a Consistent Light Attention (CLA) module within the self-attention layers of the image relighting model. As shown in Fig. 2, by incorporating additional temporally averaged features into the attention computation, CLA facilitates cross-frame interactions, producing a structurally stable lighting source. To further enhance the appearance stability across frames, we utilize the motion priors of the video diffusion model with a novel Progressive Light Fusion (PLF) strategy. Adhering to the physical principles of light transport, PLF progressively employs linear blending to integrate relighted appearances from the CLA into each original denoising target, which gradually guides the video denoising process toward the desired relighting direction. Finally, Light-A-Video serves as a complete end-to-end pipeline, effectively achieving smooth and consistent video relighting. As a training-free approach, Light-A-Video is not restricted to specific video diffusion models, making it highly compatible with a range of popular video generation backbones, including UNet-based and DiT-based models such as AnimateDiff [13], CogVideoX [60] and LTX-Video [15]. Our contributions are summarized as follows:
  • We present Light-A-Video, the first training-free video relighting model, which leverages the capabilities of image relighting models to generate high-quality, temporally consistent relighted videos.
  • An end-to-end pipeline is proposed with two key designs: the Consistent Light Attention, which enhances the stability of lighting source across frames, and the Progressive Light Fusion, which gradually injects lighting information to ensure temporal consistency for video appearance.
  • Experiments show the effectiveness of our method across multiple settings. It not only supports relighting of the entire input video, but also supports relighting the input foreground sequences with background generation.
2 Related Work
2.1 Video Diffusion Models.
Video diffusion models [3, 5, 6, 13, 51, 54, 17, 60, 2, 62, 4] aim to synthesize temporally consistent image frames based on provided conditions, such as a text prompt or an image prompt. In the realm of text-to-video (T2V) generation, the majority of methods [51, 13, 5, 6, 3, 62] train additional motion modeling modules from existing text-to-image architectures to model the correlation between video frames, while others [17, 60] train from scratch to learn video priors. For image-to-video (I2V) tasks that enhance still images with reasonable motions, a line of research [57, 7] proposes novel frameworks dedicated to image animation. Some approaches [12, 64, 14] serve as plug-to-play adapters. Stable Video Diffusion [2] fine-tune pre-trained T2V models for I2V generation, achieving impressive performance. Numerous works [34, 30, 27] focus on controllable generation, providing more controllability for users. Video diffusion models, due to their inherent video priors, are capable of synthesizing smooth and consistent video frames that are both content-rich and temporally harmonious.
2.2 Learning-based Illumination Control.
Over the past few years, a variety of lighting control techniques [48, 33, 35] for 2D and 3D visual content based on deep neural networks have been proposed, especially in the field of portrait lighting [47, 1, 44, 46, 23], along with a range of baselines [66, 45, 18, 55, 67] aimed at improving the effectiveness, accuracy, and theoretical foundation of illumination modeling. Recently, owing to the rapid development of diffusion-based generative models, a number of lighting control methods [40, 61, 10, 20] utilizing diffusion models have also been introduced. Relightful Harmonization [40] focuses on harmonizing sophisticated lighting effects for the foreground portrait conditioning on a given background image. SwitchLight [23] suggests training a physically co-designed framework for human portrait relighting. IC-Light [63] is a state-of-the-art approach for image relighting. LumiSculpt [65] enables consistent lighting control in T2V generation models for the first time. However, in the domain of video lighting, the aforementioned approaches fail to simultaneously ensure precise lighting control and exceptional visual quality. This work incorporates a pre-trained image lighting control model into the denoising process of a T2V model through progressive guidance, leveraging the latter’s video priors to facilitate the smooth transfer of image lighting control knowledge, thereby enabling accurate and harmonized control of video lighting.
2.3 Video Editing with Diffusion Models.
In recent years, diffusion-based video editing has undergone significant advancements. Some researches [29, 53, 56, 30, 32] adopt pretrained text-to-image (T2I) backbones for video editing. Another line of approaches [59, 58, 8, 19] leverages pre-trained optical flow models to enhance the temporal consistency of output video. Numerous studies [38, 11, 21] have concentrated on exploring zero-shot video editing approaches. COVE [52] leverages the inherent diffusion feature correspondence proposed by DIFT [49] to achieve consistent video editing. SDEdit [31] utilizes the intrinsic capability of diffusion models to refine details based on a given layout, enabling efficient editing for both image and video. Despite the remarkable performance of existing video editing techniques in various settings, there remains a lack of approaches specifically designed for controlling the lighting of videos.
3 Preliminary
Figure 3: The pipeline of Light-A-Video. A source video is first noised and processed through the VDM for denoising across $T_{m}$ steps. At each step, the predicted noise-free component with details compensation serves as the Consistent Target $\mathbf{z}^{v}_{0\leftarrow t}$, inherently representing the VDM’s denoising direction. Consistent Light Attention infuses $\mathbf{z}^{v}_{0\leftarrow t}$ with unique lighting information, transforming it into the Relight Target $\mathbf{z}^{r}_{0\leftarrow t}$. The Progressive Light Fusion strategy then merges two targets to form the Fusion Target $\tilde{\mathbf{z}}_{0\leftarrow t}$, which provides a refined direction for the current step. The bottom-right part illustrates the iterative evolution of $\mathbf{z}^{v}_{0\leftarrow t}$.
3.1 Diffusion Model
Given an image $\mathbf{x}_{0}$ that follows the real-world data distribution, we first encode $\mathbf{x}_{0}$ into latent space $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})$ using a pretrained autoencoder $\{{\mathcal{E}(\cdot)},{\mathcal{D}(\cdot)}\}$. The forward diffusion process is a $T$ steps Markov chain [16], corresponding to the iterative introduction of Gaussian noise $\epsilon$, which can be expressed as:
$\mathbf{z}_{t}=\sqrt{1-\beta_{t}}\mathbf{z}_{t-1}+\sqrt{\beta_{t}}\epsilon,$ (1)
where $\beta_{t}\in(0,1)$ determines the amount of Gaussian noise introduced at time step $t$. Mathematically, the above cumulative noise adding has the following closed-form:
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (2)
where $\bar{\alpha}_{t}={\textstyle\prod_{1}^{t}}(1-\beta_{t})$.
For numerical stability, $\mathbf{v}$-prediction [42] approach is employed, where the diffusion model outputs a predicted velocity $\mathbf{v}$ to represent the denoising direction. Defined as:
$\mathbf{v}=\sqrt{\bar{\alpha}_{t}}\epsilon-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}_{0}$ (3)
During inference, the noise-free component $\hat{\mathbf{z}}_{0\leftarrow t}$ can be recovered from the model’s output $\mathbf{v}_{t}$ as follows:
$\hat{\mathbf{z}}_{0\leftarrow t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{v}_{t}.$ (4)
$\hat{\mathbf{z}}_{0\leftarrow t}$ represents the denoising target based on $\mathbf{v}_{t}$.
3.2 Light Transport
Light transport theory [9, 63] demonstrates that arbitrary image appearance $\mathbf{I}_{L}$ can be decomposed by the product of a light transport matrix $\mathbf{T}$ and environment illumination $L$, which can be expressed as:
$\mathbf{I}_{L}=\mathbf{T}L,$ (5)
where $\mathbf{T}$ is a single matrix for linear light transform [9] and $L$ denotes variable environment illumination. Given the linearity of $\mathbf{T}$, the merging between environment illumination $L$ is equal to the fusion of image appearance $\mathbf{I}_{L}$, i.e.,
$\mathbf{I}_{L_{1}+L_{2}}=\mathbf{T}(L_{1}+L_{2})=\mathbf{I}_{L_{1}}+\mathbf{I}_{L_{2}}.$ (6)
Such characteristic suggests the feasibility of lighting control by indirectly constraining image appearance, i.e., the consistent image light constraint in IC-Light [63].
4 Light-A-Video
In Section 4.1, we define the objectives of the video relighting task. In Section 4.2, we reveal that per-frame image relighting for video sequence suffers from lighting source inconsistency and accordingly propose Consistent Lighting Attention (CLA) module for enhanced lighting stability in per-frame image relighting results. In Section 4.3, we provide a detailed description of the proposed Light-A-Video pipeline, which achieves stable lighting source generation and temporally consistent relighted video through Progressive Light Fusion (PLF) strategy.
4.1 Problem Formulation
Given a source video and a lighting condition $c$, the objective of video relighting is to render the source video into the relighted video that aligns with the given condition $c$. Unlike image relighting solely concentrates on appearance, video relighting raises extra challenges in maintaining temporal consistency and motion preservation, necessitating high-quality visual coherence across frames.
4.2 Consistent Light Attention
Given the achievement in image relighting model [63], a straightforward approach for video relighting is to directly perform frame-by-frame image relighting under the same lighting condition. However, as illustrated in Fig. 2, this naive method fails to maintain appearance coherence across frames, resulting in frequent flickering of the generated light source and inconsistent temporal illumination.
To improve inter-frame information integration and generate a stable light source, we propose a Consistent Light Attention (CLA) module. Specifically, for each self-attention layer in the IC-Light model, a video feature map $\mathbf{h}\in\mathbb{R}^{(b\times f)\times(h\times w)\times c}$ serves as the input, where $b$ is the batch size and $f$ is the number of video frames, $h$ and $w$ denote the height and width of the feature map, with $h\times w$ representing the number of tokens for attention computation. With linearly projections, $\mathbf{h}$ is projected into query, key and value features $Q,K,V\in\mathbb{R}^{(b\times f)\times(h\times w)\times c}$. The attention computation is defined as follows:
$\operatorname{Self-Attn}(Q,K,V)=\operatorname{Softmax}\left(\frac{QK^{T}}{\sqrt{c}}\right)V$
Note that the naive method treats the frame dimension as the batch size, performing self-attention frame by frame with the image relighting model, which results in each frame attending only to its own features. For our CLA module as shown in Fig. 3, a dual-stream attention fusion strategy is applied. Given the input $\mathbf{h}$, the original stream directly feeds the feature map into the attention module to compute frame-by-frame attention, resulting in the output $\mathbf{h}^{\prime}_{1}$. The average stream first reshapes $\mathbf{h}$ into $\mathbb{R}^{b\times f\times(h\times w)\times c}$, averages it along the temporal dimension, then expands it $f$ times to obtain $\mathbf{\bar{h}}$. Specifically, the average stream mitigates high-frequency temporal fluctuations, thereby facilitating the generation of a stable background light source across frames. Meanwhile, the original stream retains the original high-frequency details, thereby compensating for the detail loss incurred by the averaging process. Then, $\mathbf{\bar{h}}$ is input into the self-attention module and the output is $\mathbf{\bar{h}}^{\prime}_{2}$. The final output $\mathbf{h}_{o}^{\prime}$ of the CLA module is a weighted average between two streams, with the trade-off parameter $\gamma$,
$\mathbf{h}_{o}^{\prime}=(1-\gamma)\mathbf{h}^{\prime}_{1}+\gamma\mathbf{\bar{h}}^{\prime}_{2}$ (7)
With the help of CLA, the result can capture global context across the entire video, and generated a more stable relighting result as shown in Fig. 2.
4.3 Progressive Light Fusion
CLA module improves cross-frame consistency but lacks pixel-level constraints, leading to inconsistencies in appearance details. To address this, we leverage motion priors in Video Diffusion Model (VDM), which are trained on large-scale video datasets and use temporal attention module to ensure consistent motion and lighting changes. The novelty of our Light-A-Video lies in progressively injecting the relighting results as guidance into the denoising process.
Figure 4: Visualization of the PLF Strategy. During the denoising process of the VDM, the PLF strategy progressively replaces the original Consistent Target $P^{t}_{c}$ with the Fusion Target $P_{f}^{t}$, guiding the denoising direction from $\mathbf{z}^{v}_{0\leftarrow t}$ to $\tilde{\mathbf{v}}_{t}$.
In the pipeline as shown in Fig 3, a source video is first encoded into latent space, and then add $T_{m}$ step noise to acquire the noisy latent $\mathbf{z}_{m}$. At each denoising step $t$, the noise-free component $\hat{\mathbf{z}}_{0\leftarrow t}$ in Eq. 4 is predicted, which serves as the denoising target for the current step. Prior work demonstrated the potential of applying tailored manipulation in denoising targets for guided generation, with significant achievements observed in high-resolution image synthesis [24] and text-based image editing [41].
Driven by the motion priors in the VDM, the denoising process encourage $\hat{\mathbf{z}}_{0\leftarrow t}$ to be temporally consistent. Thus, we define this target as the video Consistent Target $\mathbf{z}^{v}_{0\leftarrow t}$ with environment illumination $L^{v}_{t}$. However, discrepancies still exist between the predicted $\mathbf{z}^{v}_{0\leftarrow t}$ and the original video. To address this issue, we incorporate the differences as details compensation into the Consistent Target at each step, thereby enhancing detail consistency between the relighted video and the original video. Then, $\mathbf{z}^{v}_{0\leftarrow t}$ is sent into the CLA module to obtain the relighted latent, which serves as the Relight Target $\mathbf{z}^{r}_{0\leftarrow t}$ with the illumination $L^{r}_{t}$ for the $t$-th denoising step. Aligning with the light transport theory in Section 3.2, a pre-trained VAE $\{{\mathcal{E}(\cdot)},{\mathcal{D}(\cdot)}\}$ is used to decode the two targets into pixel level, yielding the image appearances $\mathbf{I}^{v}_{t}=\mathcal{D}(\mathbf{z}^{v}_{0\leftarrow t})$ and $\mathbf{I}^{r}_{t}=\mathcal{D}(\mathbf{z}^{r}_{0\leftarrow t})$, respectively. Refer to Eq. 6, the fusing appearance $\mathbf{I}^{f}_{t}$ can be formulated as,
$\mathbf{I}^{f}_{t}=\mathbf{T}(L^{v}_{t}+L^{r}_{t})=\mathbf{I}^{v}_{t}+\mathbf{I}^{r}_{t}.$ (8)
We observe that directly using encoded latent $\mathcal{E}(\mathbf{I}^{f}_{t})$ as the new target at each step results in suboptimal performance. The reason is that the gap between the two targets is excessively large, surpassing the refinement capability of the VDM and causing visible temporal lighting jitter. To mitigate this gap, a progressive lighting fusion strategy is proposed. Specifically, a fusion weight $\lambda_{t}$ is introduced, which decreases as denoising progresses, gradually reducing the influence of the relight target. The progressive light fusion appearance is defined as $\mathbf{I}^{p}_{t}$,
$\mathbf{I}^{p}_{t}=\mathbf{I}^{v}_{t}+\lambda_{t}(\mathbf{I}^{r}_{t}-\mathbf{I}^{v}_{t}).$ (9)
The encoded latent $\tilde{\mathbf{z}}_{0\leftarrow t}=\mathcal{E}(\mathbf{I}^{p}_{t})$ is utilized as the Fusion Target for step $t$, replacing the original $\hat{\mathbf{z}}_{0\leftarrow t}$. Based on the fusion target, the less noisy latent $\mathbf{z}_{t-1}$ can be computed with DDIM scheduler with $v$-prediction:
$a_{t}=\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}},\\ b_{t}=\sqrt{\bar{\alpha}_{t-1}}-\sqrt{\bar{\alpha}_{t}}a_{t}$ (10)
$\mathbf{z}_{t-1}=a_{t}\mathbf{z}_{t}+b_{t}\tilde{\mathbf{z}}_{0\leftarrow t}$ (11)
From Eq. 4, the fusion target $\tilde{\mathbf{z}}_{0\leftarrow t}$ determines a new denoising direction, denoted as $\tilde{\mathbf{v}}_{t}$,
$\tilde{\mathbf{v}}_{t}=\frac{\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{t}-\tilde{\mathbf{z}}_{0\leftarrow t}}{\sqrt{1-\bar{\alpha}_{t}}},$ (12)
which means our PLF strategy essentially refines $\mathbf{v}_{t}$ iteratively and guides the denoising process towards the relighting direction shown in Fig 4. Notably, other schedulers (e.g., Euler Scheduler [22], Flow Matching [28], etc.) with predicted noise-free components for each step are also applicable. Finally, as the denoising progresses, we achieve a smooth and consistent injection of illumination, ensuring coherent video relighting.
5 Experiments
Figure 5: Qualitative comparison of baseline methods. Given a source video and guidance text prompt, we compare the performance of Light-A-Video with other video editing methods. VDM used: AnimateDiff (Left), CogVideoX (Right).
In this section, we provide comprehensive qualitative and quantitative evaluations to highlight the efficacy and generality of our method for zero-shot video illumination control. Furthermore, we conduct ablation studies to analyze the contributions of critical model components and provide additional results on text-conditioned video illumination modifying with background generation.
5.1 Experimental Details
Baselines. In the absence of established video relighting methods, we adopt the state-of-the-art image relighting technique to perform frame-by-frame relighting on videos as a baseline. To verify the temporal smoothing effect of illumination using a VDM, we add 20% and 60% noise to the results obtained from IC-Light [63]. Subsequently, we apply SDEdit [31] on VDM for denoising, resulting in two sets of results named IC-Light + SDEdit-0.2 and IC-Light + SDEdit-0.6, respectively. Finally, AnyV2V [26] is utilized as another baseline to verify whether the temporal content migration method can be effectively applied to video relighting tasks. This baseline names IC-Light + AnyV2V, where the first frame of the IC-Light result and the source video are used as conditional inputs.
Evaluation metrics. We report three widely adopted metrics for quantitative evaluation. First, to evaluate the temporal consistency of the generated video, the average CLIP [39] score across consecutive video frames is utilized. Second, RAFT [50] is used to estimate the optical flow for each baseline video, and assess the motion preservation of each method by calculating the optical flow similarity with the source video. Third, to evaluate the quality of image relighting for each method, a video test dataset is collected, and the FID [43] score is calculated between each method’s results And the frame-by-frame IC-Light results, serving as the relight quality evaluation metric.
Datasets. We constructed a video test dataset consisting of 73 videos. The majority of these videos were selected from the DAVIS [37] public dataset, which contains a diverse collection of semantically rich videos with pronounced motion. Additionally, some videos were collected from Pixabay [36], featuring high-quality videos with significant motion. All quantitative metrics are evaluated on our collected dataset. For each video, two lighting prompts are applied, and three lighting directions are randomly chosen.
Implementation details. If not specified, the default models employed for image relighting and VDM in the subsequent experiments are IC-Light [63] and AnimateDiff [13] motion-adapter-v3, respectively. In the IC-Light model, the lighting conditions $c$ for image relighting are derived from two components. First, a text prompt that describes the characteristics of the light source (e.g., neon light, sunshine, etc.). Second, a lighting map is utilized to represent the light intensity across the scene. This lighting map is then encoded by a VAE and serves as the initial latent for the denoising process. During the inference stage, the source video is added with 50% noise. Then, the VDM employs a denoising process with $T_{m}=25$ steps to progressively fusion the relight target. For the parameters in our Light-A-Video model, the $\gamma=0.5$ in the CLA module is used to balance the original attention hidden state and the cross-frame averaged hidden state. In our PLF strategy, the fusion weight $\lambda_{t}$ is decreases as denoising progresses, so we set $\lambda=1-t/T_{m}$. All experiments were conducted on a single A100 GPU, with a memory consumption of approximately 23GB for a 16-frame relighted video generation.
Figure 6: Text-conditioned video illumination modifying with background generation. Given a video foreground sequence and a text description of the target illumination, our method synthesizes suitable backgrounds and harmonious illumination.
Figure 7: Ablation Study. We present the video relighted with background generation results by removing the CLA module or the PLF strategy.
5.2 Qualitative Results
As shown in Fig. 5, the frame-by-frame IC-Light method ensures single-frame quality. However, the absence of consistency design and VDM temporal priors results in significant flickering of the light source and overall appearance. By introducing VDM priors, IC-Light + SDEdit-0.2 maintains content consistent with the source video, but still exhibits noticeable relight appearance jitter. IC-Light + SDEdit-0.6 further enhances temporal smoothness, yet object identity shifts occur. AnyV2V can transfer the appearance of the first relight frame to subsequent frames, but this type of temporal content migration method inherently performs only pixel-level migration without perceiving the given light source, leading to unreasonable changes in object illumination. In contrast, Light-A-Video achieves high-quality video relighting, demonstrating strong temporal consistency and high fidelity to the light source.
5.3 Video Relighting with Background Generation
As depicted in Fig. 6, Light-A-Video not only supports the direct relighting of the input source video but also accepts a video foreground sequence and a text prompt provided by users as input, generating a corresponding video background and illumination that aligned with the prompt descriptions. Specifically, we first process the input foreground sequence with IC-Light for frame-by-frame relighting, while completely noising the background to serve as the initialization video for VDM. From step $T$ to $T_{m}$, the background is gradually denoised, leveraging VDM’s inpainting capability to generate the background. Then, from step $T_{m}$ to 0, CLA and PLF are introduced to achieve a temporally consistent relighting appearance of the video. The results in Fig. 6 demonstrate that our pipeline can achieve high-quality video relighting results with consistent background generation.
5.4 Quantitative Evaluation
The quantitative comparison of our method with various baselines is presented in Tab. 1. As shown, we introduced an additional baseline SDEdit-0.6, which denoises the source video with 60% noise. SDEdit-0.6 results in a high temporal Clip Score demonstrating relying solely on the relight prompt is insufficient for generating videos with the corresponding lighting characteristics using VDM. When the frame-by-frame IC-Light results are used as the initialization video and then noised, IC-Light + SDEdit-0.6 exhibits an increased FID score, indicating a trade-off with temporal stability. Similarly, IC-Light + SDEdit-0.2, which incorporates 20% noise, maintains a high consistency in lighting effects with frame-by-frame IC-Light, but it also suffers from unacceptable temporal flickering. For AnyV2V, a temporal content migration method, the first frame’s appearance is consistent with the IC-Light results. Yet, its inability to perceive the light source, coupled with the model’s inherent quality degradation in subsequent frames, results in a low motion preservation score. In contrast, our Light-A-Video method achieves a low FID score while maintaining a high level of temporal consistency, demonstrates the effectiveness in both relighted image quality and temporal stability.
Evaluation Metric (a) Image Quality (b) Temporal Consistency
FID Score($\downarrow$) CLIP Score($\uparrow$) Motion Preservation($\downarrow$)
IC-Light [63] / 0.9040 5.969
SDEdit-0.6 [31] 64.03 0.9677 4.204
IC-Light + SDEdit-0.2 13.79 0.9199 5.959
IC-Light + SDEdit-0.6 62.61 0.9483 7.544
IC-Light + AnyV2V [26] 32.73 0.9436 8.854
Light-A-Video (Ours) 29.63 0.9655 2.493
Table 1: Quantitative comparison of baseline methods. The best result is highlighted in bold, while the second-best result is underlined.
5.5 Ablation Study
We conduct an ablation study to understand the importance of our CLA and PLF modules. As shown in Fig. 7, for the video relighting with background generation task, frame-by-frame IC-Light ensures high single-frame relighting quality but lacks temporal consistency control, leading to lighting source inconsistency and relighted appearance inconsistency. With the help of CLA, cross-frame information exchange stabilizes the generation of background lighting sources. Furthermore, by introducing VDM motion priors and using PLF’s strategy to progressively fuse relight targets into the original denoising target, Light-A-Video achieves temporally smooth relighting of the video. The overall video quality is also significantly improved with the help of VDM priors.
5.6 Limitation and Future Work.
Despite achieving impressive results, the performance of our training-free method is limited by the capabilities of both the image-relighting model and the VDM. While Light-A-Video excels at ensuring stable lighting and temporal consistency, the CLA module, which stabilizes the background lighting, is less effective for modeling dynamic lighting changes. To address this limitation, future work will focus on developing methods that can handle dynamic lighting conditions more effectively.
6 Conclusion
In summary, our paper introduces Light-A-Video, a training-free method that leverages the capabilities of state-of-the-art image relighting models to achieve temporally consistent video relighting. By integrating the Consistent Light Attention to stabilize lighting source generation and employ the Progressive Light Fusion strategy for smooth appearance transitions. Light-A-Video significantly enhances the temporal coherence of relighted videos while preserving the high-quality relighting of images. Experiments demonstrate the effective application of advanced image-relighting techniques in both video sequence relighting and foreground relighting with background generation.